# DreamRunner: Fine-Grained Storytelling Video Generation with Retrieval-Augmented Motion Adaptation

### v2

@@ Appendix A Ablations of RAG pipeline
We provide additional experiments in Section 3.2 to demonstrate the effectiveness of our data processing approach within the retrieval pipeline. Table 1 primarily evaluates the impact of the maximal number of retrieved videos and the use of CLIP and ViCLIP for filtering. For efficiency, we evaluate a subset of eight motions selected from the full pool of 64 motions. The results indicate that, compared to the CogVideoX zero-shot baseline (Row 1), retrieving videos without any filtering (Row 2) improves performance, even though BM-25 [robertson2009probabilistic] retrieval introduces some noise. This highlights the importance of retrieval itself. Furthermore, adding CLIP and ViCLIP as filters (Row 4) further enhances performance, showcasing the benefit of using semantically aligned videos for improved motion learning. Additionally, retrieving a sufficient number of videos is critical, as evidenced by Row 3, where limiting retrieval to a maximum of three videos results in poorer performance compared to retrieving 20 videos (Row 4).
@@
Appendix B Evaluation Metrics
@@
We provide detailed evaluation metrics for evaluating generated storytelling videos in multiple dimensions in Section 4.1.
@@
  • Similarity to Reference Images: We assess the alignment between the generated videos and reference images using Image-Image CLIP [hessel2021clipscore] and DINO [oquab2023dinov2] scores. The evaluation is conducted by averaging the CLIP/DINO similarity scores with each frame of the generated video and the reference images, with CLIP-L14 (336px) [hessel2021clipscore] and DINOv2-B16 [oquab2023dinov2] as image encoders.
@@
  • Similarity to Full Narration per scene: We measure the alignment between the full narration and the generated videos for each scene using Image-Text CLIP [hessel2021clipscore] and Video-Text ViCLIP [wang2023internvid] scores. For the CLIP score, we uniformly sample eight frames from the single-scene video and compute the average score between each frame and the full narration. For the ViCLIP score, we directly use the alignment score between the video and the narration. Per-scene scores are averaged to obtain the overall score for the multi-scene storytelling video.
@@
  • Fine-Grained Text Alignment per Scene: We also assess fine-grained alignment to textual descriptions using Image-Text CLIP [hessel2021clipscore] and Video-Text ViCLIP [wang2023internvid] scores. In our story, each narration contains two motions, which we decouple into two consecutive single-motion descriptions using an LLM [openai_gpt4o]. For each generated single-scene video, we divide it into two segments at the temporal midpoint. We then compute the CLIP/ViCLIP scores between the first segment and the first description and between the second segment and the second description. The average of these two scores provides the per-scene score, and the per-scene scores are further averaged to compute the overall score for the storytelling video.
@@
  • Transition: We assess whether the single-scene video achieves smooth transitions between two motions. To evaluate this, we uniformly sample four frames from the video per scene and calculate the average DINO similarity between adjacent frames. A higher transition score indicates smoother transitions, as it reflects minimal changes in the background across frames.
Appendix C Compositional T2V Examples
@@
For action binding, as shown in Figure 1, DreamRunner generates distinct motions for two objects (e.g., a wolf howling into a microphone and a fox playing the drums), ensuring the actions are correctly bound to their respective objects without interference.
For consistent attribute binding, as illustrated in Figure 2, our approach maintains separate attributes for different objects (e.g., a spherical globe and a cube-shaped clock) without any overlap or inconsistency.
For dynamic attribute changes, as shown in Figure 3, DreamRunner naturally transitions object attributes over time (e.g., metal gradually rusting throughout the video).
For motion control, as depicted in Figure 4, DreamRunner successfully directs object movements in different trajectories (e.g., a kite moving from left to right and a cyclist traveling from right to left).
For object interactions, as illustrated in Figure 5, interactions are accurately modeled, adhering to physical rules (e.g., pottery shattering upon hitting the floor).
@@
Appendix D Characters.
We provide character examples in Figure 7, where the first four are generated using FLUX [flux], and the others are collected from existing customization datasets (CustomConcept101 [kumari2023multi] and Dreambooth [ruiz2023dreambooth]).
Appendix E Single-Character Examples
@@
Appendix F Multi-Character Examples
@@
Appendix G Region-Based 3D Attention Masks
Figure 12 shows and example of the attention Mask of our Spatial-Temporal Region-based 3D Attention introduced in Section LABEL:sec:subsec:attn. Different text colors represent different conditions, while the white region indicates the masked areas. Note that for simplicity, we reduce each condition to two words, each frame to three segments, and display only three conditions and two frames in the figure. In practice, conditions can be longer and more numerous, frames can have more segments, and there are 12 latent frames in total.
@@ Appendix H LLM Prompts
Figure 1: Qualitative results of DreamRunner generated with prompts characterizing action binding.
Figure 2: Qualitative results of DreamRunner generated with prompts characterizing consistent attribute binding.
Figure 3: Qualitative results of DreamRunner generated with prompts characterizing dynamic attribute binding.
Figure 4: Qualitative results of DreamRunner generated with prompts characterizing motion binding.
Figure 5: Qualitative results of DreamRunner generated with prompts characterizing object interactions.
Figure 6: Qualitative results of DreamRunner generated with prompts characterizing spatial relationships.
@@
Figure 12: Visualization of spatial-temporal region-based 3D attention masks. Different text colors represent different conditions, while the white region indicates the masked areas. For simplicity, we reduce each condition to two words, each frame to three segments, and display only three conditions and two frames in the figure. In practice, conditions can be longer and more numerous, frames can have more segments, and there are 12 latent frames in total.

### v3

# DreamRunner: Fine-Grained Compositional Story-to-Video Generation with Retrieval-Augmented Motion Adaptation

@@ B.1 Ablations of RAG pipeline
We provide additional experiments in Section 3.2 to demonstrate the effectiveness of our data processing approach within the retrieval pipeline. Table 5 primarily evaluates the impact of the maximal number of retrieved videos and the use of CLIP and ViCLIP for filtering. For efficiency, we evaluate a subset of eight motions selected from the full pool of 64 motions. The results indicate that, compared to the CogVideoX zero-shot baseline (Row 1), retrieving videos without any filtering (Row 2) improves performance, even though BM-25 [49] retrieval introduces some noise. This highlights the importance of retrieval itself. Furthermore, adding CLIP and ViCLIP as filters (Row 4) further enhances performance, showcasing the benefit of using semantically aligned videos for improved motion learning. Additionally, retrieving a sufficient number of videos is critical, as evidenced by Row 3, where limiting retrieval to a maximum of three videos results in poorer performance compared to retrieving 20 videos (Row 4).
@@
B.2 Ablations of Odd-Even Layer Separation for Motion Prior Learning
During motion prior learning, since our backbone does not explicitly separate spatial and temporal modeling, we manually designate even layers as spatial and odd layers as temporal to decouple the learning of spatial and temporal LoRAs. We hypothesize that this design helps learn appearance-debiased, motion-focused priors, while the interleaved structure allows adaptation across low-level structures and high-level semantics, as both spatial and temporal LoRAs are injected throughout the model. We validate this hypothesis in Table 6. Specifically, removing spatial LoRAs (row 1) degrades performance compared to row 3, indicating that separating spatial and temporal LoRAs enhances motion-focused learning. Additionally, the Half-Half approach (injecting spatial LoRAs into the first half of layers and temporal LoRAs into the second half, row 2) underperforms compared to interleaved injection, highlighting the importance of multi-level learning across the entire model.
  Method CLIP ViCLIP
No Appearance-Debiased 24.4 22.5
Half-Half Injection 24.9 23.1
Interleaved Injection (Ours) 25.5 23.7
Table 6: LoRA injection ablations.
Appendix C Evaluation Metrics
@@
We provide detailed evaluation metrics for assessing generated storytelling videos across multiple dimensions in Section 4.1.
@@
  • Similarity to Reference Images: We assess the alignment between the generated videos and reference images using Image-Image CLIP [19] and DINO [42] scores. The evaluation is conducted by averaging the CLIP/DINO similarity scores between each frame of the generated video and the reference images, with CLIP-L14 (336px) [19] and DINOv2-B16 [42] as image encoders.
@@
  • Similarity to Full Narration per Scene: We measure the alignment between the full narration and the generated videos for each scene using Image-Text CLIP [19] and Video-Text ViCLIP [58] scores. For the CLIP score, we uniformly sample eight frames from the single-scene video and compute the average score between each frame and the full narration. For the ViCLIP score, we directly use the alignment score between the video and the narration. Per-scene scores are averaged to obtain the overall score for the multi-scene storytelling video.
@@
  • Fine-Grained Text Alignment per Scene: We also assess fine-grained alignment to textual descriptions using Image-Text CLIP [19] and Video-Text ViCLIP [58] scores. In our story, each narration contains two motions, which we decouple into two consecutive single-motion descriptions using an LLM [41]. For each generated single-scene video, we divide it into two segments at the temporal midpoint. We then compute the CLIP/ViCLIP scores between the first segment and the first description and between the second segment and the second description. The average of these two scores provides the per-scene score, and the per-scene scores are further averaged to compute the overall score for the storytelling video.
@@
  • Transition Quality: We assess whether the single-scene video achieves smooth transitions between two motions. To evaluate this, we uniformly sample four frames from the video per scene and calculate the average DINO similarity between adjacent frames. A higher transition score indicates smoother transitions, as it reflects minimal changes in the background across frames.
  • Motion Smoothness: Following [23], we utilize motion priors from a video frame interpolation model [34] to quantify the smoothness of motion in generated videos (see Supplementary for details), specifically, given one video, we drop the odd-number frames and then use the interpolation model to infer the dropped frames. Then we compute the Mean Absolute Error between the estimated frames and the dropped frames.
  • Aesthetic Quality: We evaluate the artistic and visual appeal of each generated video frame using the LAION aesthetic predictor [51] frame-by-frame. This metric reflects aesthetic aspects such as composition, color richness and harmony, photorealism, naturalness, and overall artistic quality.
  • Imaging Quality: Imaging quality measures distortions such as over-exposure, noise, and blur in generated frames. We evaluate this using the MUSIQ [27] image quality predictor frame-by-frame, which is trained on the SPAQ dataset [12].
Appendix D Compositional T2V Examples
@@
For action binding, as shown in Figure 5, DreamRunner generates distinct motions for two objects (e.g., A monkey in a pilot jacket and a parrot with aviator goggles ready for flight), ensuring the actions are correctly bound to their respective objects without interference.
For consistent attribute binding, as illustrated in Figure 6, our approach maintains separate attributes for different objects (e.g., Rectangular briefcase swinging near a hexagonal fountain) without any overlap or inconsistency.
For dynamic attribute changes, as shown in Figure 7, DreamRunner naturally transitions object attributes over time (e.g., A timelapse of a piece of bread initially fresh, then growing moldy).
For motion control, as depicted in Figure 8, DreamRunner successfully directs object movements in different trajectories (e.g., A squirrel climbs upward on a tree and A drone is gradually descending to the ground in a park).
For object interactions, as illustrated in Figure 9, interactions are accurately modeled, adhering to the real world rules (e.g., Cat’s paw presses on a soft pillow).
@@
Appendix E Characters.
We provide character examples in Figure 11, where the first four are generated using FLUX [4], and the others are collected from existing customization datasets (CustomConcept101 [30] and Dreambooth [50]).
Appendix F Single-Character Examples
@@
Appendix G Multi-Character Examples
@@ Appendix H LLM Prompts
Figure 5: Qualitative results of DreamRunner generated with prompts characterizing action binding. SR3A denotes our spatial-temporal region-based attention module.
Figure 6: Qualitative results of DreamRunner generated with prompts characterizing consistent attribute binding. SR3A denotes our spatial-temporal region-based attention module.
Figure 7: Qualitative results of DreamRunner generated with prompts characterizing dynamic attribute binding. SR3A denotes our spatial-temporal region-based attention module.
Figure 8: Qualitative results of DreamRunner generated with prompts characterizing motion binding. SR3A denotes our spatial-temporal region-based attention module.
Figure 9: Qualitative results of DreamRunner generated with prompts characterizing object interactions. SR3A denotes our spatial-temporal region-based attention module.
Figure 10: Qualitative results of DreamRunner generated with prompts characterizing spatial relationships. SR3A denotes our spatial-temporal region-based attention module.